Calculate the amount of heat supplied to a carnot cycle working between 368K and 288K if maximum obtain is 895J.
1
Expert's answer
2020-05-03T15:23:06-0400

maximum efficiency in carnot cycle ="\frac{T_2-T_1}{T_2}=\frac{368-288}{368}=0.217"


Heat obtained = efficiency"\times"heat supplied

895=0.217*heat supplied

heat supplied = 4124 J